# Merging Rate of Opinions via Optimal Transport on Random Measures

### v2

@@ 1 Introduction
$X_{1},\dots,X_{n}|\tilde{\mu}\stackrel{{\scriptstyle\mbox{\scriptsize{\rm iid}}}}{{\sim}}t(\tilde{\mu}),$ (1)
@@
Before moving to the analysis of (2), one needs to investigate the identifiability of the random parameter $\tilde{\mu}$ in (1), that is, which random measures $\tilde{\mu}$ induce the same $t(\tilde{\mu})$. That is crucial to avoid outputting a strict positive distance even though the laws of the observation are the same, that is, to erroneously detect a lack of merging of opinion. Identifiability depends on the choice of transformation $t$ and the class of distributions for $\tilde{\mu}$: we focus on a specific setting, namely, normalization of CRMs. The task then amounts to understand which random measures induce the same normalization, divide the random measures in equivalence classes accordingly, and choose a representative for each equivalence class. This will lead to the definition of scaled CRMs $\tilde{\mu}_{\mathcal{S}}$, such that $\mathbb{E}(\tilde{\mu}_{\mathcal{S}}(\mathbb{X}))=1$, and their generalizations to Cox CRMs.
@@
This sets the ground for a principled framework for studying the merging rate of opinions in terms of the distance between (the representatives of) the posteriors, as in (2). Relying on the almost-conjugacy of models (1) with respect to a CRM prior (James et al., 2009), we are able to shed new light on the use of the Dirichlet process and the normalized generalized gamma process, with unexpected results both from the asymptotic (Questions a. and b.) and finite-sample (Question c.) point of view.
@@
We first focus on the Dirichlet process (Ferguson, 1973), which has been a cornerstone in the development of BNP, allowing one to make inference on the law of the observations through closed-form expressions. We use our setup to investigate to which extent different parameters for the Dirichlet process will lead to different learning outcomes, both in a finite sample and asymptotic scenario, without putting any assumption on the “true” data-generating process. In this context, we prove that the merging will always occur at a rate 1/$n$, where $n$ is the sample size, independently of the dimension of the space $\mathbb{X}$ and model’s parameters. Moreover, we show that if the two Dirichlet processes have the same concentration parameter the merging will start immediately with the first observation, whereas if they differ it is not always the case, even if the base probabilities $P_{0}$ coincide. From a statistical perspective this finding is quite remarkable: even if two Bayesians have the same prior guess $P_{0}$ of the distribution of the data, the difference between their opinions may increase if they see a finite amount of data, according to how much confidence they are willing to put on their prior guesses.
@@
Next we move to generalized gamma CRMs (Brix, 1999), a versatile class of CRMs that recovers many well-known CRMs as special or limiting cases. This is often invoked as a generalization of the Dirichlet process that adds a discount parameter $\sigma$ to increase the flexibility of the model. Its normalization has been introduced by Lijoi et al. (2007) and used in many statistical analyses (e.g., De Blasi et al., 2009, Barrios et al., 2013, Griffin and Leisen, 2017, Caron and Fox, 2017, Lau and Cripps, 2022, Lee et al., 2023). We use our setup to investigate its merging with the Dirichlet process, again without putting any assumption on the data-generating process. We prove that the merging rate occurs at the rate $\max(n^{-1/(1+\sigma)},k/n)$, where $k$ is the number of distinct observations in the dataset. This result unravels the regimes where using a normalized generalized gamma CRM will actually guarantee a different learning outcome with respect to the Dirichlet process, with the following take-home message: either the number of observations $n$ is sufficiently small or the number of distinct values $k$ should be sufficiently large, and how large it must be depends on the choice of discount $\sigma$, namely, $k>n^{\frac{\sigma}{1+\sigma}}$.
@@
Optimal transport between random measures. The idea of endowing the space of measures, or probability measures, with a distance and then considering an optimal transport distance on laws of random measures with cost function being this distance between measures is not new. This has been done between laws of random probability measures when the distance is an optimal transport distance with recent applications in machine learning (Yurochkin et al., 2019, Dukler et al., 2019) and Bayesian statistics (Nguyen, 2016, Catalano and Lavenant, 2024). In a different direction, Huesmann (2016) extended this “OT over OT” problem to the case of random measures which are equivariant, which implies that the law of the random measures is invariant under translation. This corresponds to measures with infinite mass (because the space is unbounded), and the average transport cost is used as a measure of distance. Erbar et al. (2023) then studied the gradient flows in the geometry generated by this optimal transport problem. There have also been proposals of distances tailored for laws of point processes (see, e.g., Barbour and Brown, 1992, Schuhmacher and Xia, 2008, Decreusefond et al., 2016, and references therein), but in his work we look at random measures, not only point processes. In general terms, this forces us to consider also the distribution of the jumps of the random measures, and not only the one of their atoms.
Optimal transport on CRMs. The first proposals to define a distance between CRMs with tractable upper bounds can be found in Catalano et al. (2020, 2021), where the authors propose to use
@@
which is a distance between CRMs (Catalano et al., 2024) and leads to closed-form expressions whenever i) the CRMs do not have fixed atoms; ii) the CRMs are almost surely (a.s.) finite and homogeneous, i.e., $\nu^{i}=\rho^{i}\otimes P^{i}$ is a product measure; iii) the CRMs share the same base measure, i.e., $P^{1}=P^{2}$. These conditions are usually met in the setting of Catalano et al. (2024) but are very restrictive in our framework since i) and ii) typically hold a priori but not a posteriori, and iii) entails that the atoms of the CRM have the same distribution, which is likely to be the first feature to change when choosing a different prior. Therefore we need to define the adapted extended Wasserstein discrepancy, a new OT distance on Lévy intensities with same mean measure that considers the joint distribution of the atoms and the jumps instead. Remarkably, it coincides with (3) whenever conditions i), ii), and iii) apply (see Remark 2 below).
@@
Merging of opinions. The notion of merging of opinions was introduced and thoroughly studied in the acclaimed work of Blackwell and Dubins (1962), which holds for a very general setup beyond exchangeability. In principle their results cover model (1). However, in this setting, their assumptions can be considered restrictive for at least two reasons. First of all the merging is guaranteed for prior laws on $(X_{n})_{n\geq 1}$ that are absolutely continuous with one another. Whereas this holds in most parametric settings, it is not usually the case in nonparametric ones, where non-dominated models are the norm. As a classical example, two Dirichlet processes with different total base measures are mutually singular (Korwar and Hollander, 1973). It has also been observed by Diaconis and Freedman (1986), who avoid this condition by assessing the merging under a weaker topology. Moreover, both Blackwell and Dubins (1962) and Diaconis and Freedman (1986) assume that the data-generating mechanism coincides with one of the models under consideration, which may well not be the case in practice. Our aim is to overcome both restrictions and build a framework that does not require any assumption on the data. Indeed, we propose one of the few asymptotic analyses that does not put strong assumptions on the distribution of the data (e.g., to be a sampled from the model or i.i.d. from a true distribution): this is possible because we focus on the similarities between specific versions of the posteriors rather than on the adherence of the learning mechanism to the unknown truth.
@@
Frequentist consistency and contraction rates. In the same spirit, our framework goes beyond frequentist consistency properties in that it does not assume the data to be i.i.d. from a true distribution. However, if one were willing to make this assumption, consistency of the model with respect to each prior would entail merging of opinions but would not provide meaningful information about the convergence rate. Indeed the contraction rate of each prior to the truth could be (and usually is) considerably slower than the merging rate. For example, in this work we show that two Dirichlet processes with different parameters always merge at the rate 1/$n$, independently of the dimension of the space. This is much faster than the parametric rate of $1/\sqrt{n}$, which is the best possible one for the contraction rate, and moreover it is well known that the latter deteriorates when the dimension increases. Thus, not only our framework has the advantage of not putting assumptions on the data-generating process, it also provides new prospectives when one is willing to put them.
@@
Sensitivity analysis and robustness. To our knowledge, this work provides the first theoretical setup of global sensitivity analysis in Bayesian nonparametric non-dominated models, in the sense that it studies the performance of the model of interest with respect to any choice of prior. Previous studies on sensitivity for BNP models usually consist in varying the parameters of the priors or taking functional perturbations (Nieto-Barajas and Prünster, 2009) and providing empirical comparisons between the posteriors. In a dominated setting, more classical distances between (posterior) densities can be considered, as in Saha and Kurtek (2019). A recent work of Giordano et al. (2022) focuses on local sensitivity, which studies robustness with respect to small perturbation of the prior through differential approximations (Gustafson, 1996) and can be very useful when posterior distributions are not known in closed form. Since local sensitivity focuses on small perturbations, a good impact is always a small impact. On the contrary, our framework does not limit to small perturbations, and thus a big impact of the prior can also be regarded as a benign sign of flexibility of the model. Although our statistical analysis focuses on the latter aspect, our results may be read or adapted to treat robustness as well.
@@
Structure of the paper. In Section 2 we define the Wasserstein over BL distance between the laws of random measures. We characterize its topology and provide a general lower bound, specializing it to Poisson random measures. In Section 3 we focus on CRMs and provide a universal upper bound in terms of the adapted extended Wasserstein discrepancy between their Lévy intensities. The bound is specialized to Poisson random measures and extended to Cox CRMs with a random Lévy intensity. In Section 4 we establish the class of (Cox) CRMs that are identifiable for model (1) with respect to normalization, leading to the notion of scaled (Cox) CRMs. The adapted extended Wasserstein discrepancy is a distance on this class: in Section 5 we study its finite sample and asymptotic behaviour to investigate the merging of opinions for BNP models. All proofs are deferred to the supplementary material. Table 1 summarizes the distances that we use in this work.
@@
$\mathcal{W}_{d_{\mathbb{X}}}$ $1$-Wasserstein distance. $\mathcal{P}(\mathbb{X})$ Eq. (4)
$\mathcal{W}_{d_{\mathbb{X},t}}$ $1$-Wasserstein distance with metric $d_{\mathbb{X},\text{t}}$. $\mathcal{P}(\mathbb{X})$ Eqs. (4) and (6)
@@ 2.1 Distances between measures
[…]
where the minimum is taken over $\Pi(p^{1},p^{2})$ the set of joint probability distributions having marginals $p^{1},p^{2}$, and the equality follows by Kantorovich duality. Here $1$-Lipschitz means $|f(x)-f(y)|\leq d_{\mathbb{Y}}(x,y)$ for all $x,y\in\mathbb{Y}$. In the sequel we will use this formula for $(\mathbb{Y},d_{\mathbb{Y}})=(\mathbb{X},d_{\mathbb{X}})$, but also when $\mathbb{Y}=\mathcal{M}_{B}(\mathbb{X})$ or $\mathbb{Y}=\mathcal{M}((0,+\infty)\times\mathbb{X})$ is a space of measures.
Restricting to the case $(\mathbb{Y},d_{\mathbb{Y}})=(\mathbb{X},d_{\mathbb{X}})$, note that with $\mathcal{W}_{d_{\mathbb{X}}}$ we cannot compare measures of different mass, that is, $\mathcal{W}_{d_{\mathbb{X}}}(p^{1},p^{2})$ becomes infinite if $p^{1}(\mathbb{X})\neq p^{2}(\mathbb{X})$ as can be seen by taking constant functions $f$. A natural generalization is the Bounded Lipschitz distance $\mathrm{BL}$, which replaces the supremum over $1$-Lipschitz functions in (4) by a supremum over $1$-bounded and $1$-Lipschitz functions. Specifically we define
[…]
where $1$-bounded means $|f(x)|\leq 1$ for all $x\in\mathbb{X}$. The distance $\mathrm{BL}$ metrizes the weak convergence on $\mathcal{M}_{B}(\mathbb{X})$ (Düll et al., 2021, Theorem 1.57). Regarding the comparison between $\mathcal{W}_{d_{\mathbb{X}}}$ and $\mathrm{BL}$ on $\mathcal{P}(\mathbb{X})$, note that we always have $\mathrm{BL}\leq\mathcal{W}_{d_{\mathbb{X}}}$. However the discrepancy can be large: $\mathrm{BL}$ is always bounded by $2$, while $\mathcal{W}_{d_{\mathbb{X}}}$ can be arbitrary very large if the first moments of the measures are large. To say something finer we introduce the “truncated” distance
@@
Thus $\mathcal{W}_{d_{\mathbb{X},\text{t}}}$ and $\mathrm{BL}$ are strongly equivalent metrics when restricted to $\mathcal{P}(\mathbb{X})$ but $\mathcal{W}_{d_{\mathbb{X},\text{t}}}$ easier to interpret as it corresponds to a classical Wasserstein distance, though with respect to a modified distance.

### v3

@@ 1 Introduction
[…]
@@
[…]
@@
This sets the ground for a principled framework to study the merging rate of opinions in terms of the distance between (the representatives of) the posteriors, as in (2). Relying on the almost-conjugacy of models (1) with respect to a CRM prior (James et al., 2009), we are able to shed new light on the use of the Dirichlet process and the normalized generalized gamma process, with unexpected results both from the asymptotic (Question a. and b.) and finite-sample (Question c.) point of view.
@@
We first focus on the Dirichlet process (Ferguson, 1973), which has been a cornerstone in the development of BNP, allowing one to make inference on the law of the observations through closed-form expressions. We use our setup to investigate to which extent different parameters for the Dirichlet process will lead to different learning outcomes, both in a finite sample and asymptotic scenario, without putting any assumption on the “true” data-generating process. In this context, we prove that the merging in Wasserstein over BL will always occur at a rate faster than 1/$n$, where $n$ is the sample size, independently of the dimension of the space $\mathbb{X}$ and model’s parameters. Moreover, we show that if the two Dirichlet processes have the same concentration parameter the merging in adapted extended Wasserstein distance will start immediately with the first observation, whereas if they differ the distance may first start increasing, even if the base probabilities $P_{0}$ coincide. From a statistical perspective this finding is quite remarkable: even if two Bayesians have the same prior guess $P_{0}$ of the distribution of the data, the difference between their opinions may increase if they see a finite amount of data, according to how much confidence they are willing to put on their prior guesses.
@@
Next we move to generalized gamma CRMs (Brix, 1999), a versatile class of CRMs that recovers many well-known CRMs as special or limiting cases. This is often invoked as a generalization of the Dirichlet process that adds a discount parameter $\sigma$ to increase the flexibility of the model. Its normalization has been introduced by Lijoi et al. (2007) and used in many statistical analyses (e.g., De Blasi et al., 2009, Barrios et al., 2013, Griffin and Leisen, 2017, Caron and Fox, 2017, Lau and Cripps, 2022, Lee et al., 2023). We use our setup to investigate its merging with the Dirichlet process, again without putting any assumption on the data-generating process. We prove that the merging rate occurs at the rate $\max(n^{-1/(1+\sigma)},k/n)$, where $k$ is the number of distinct observations in the dataset. This result unravels the regimes where using a normalized generalized gamma CRM will actually guarantee a different learning outcome with respect to the Dirichlet process, with the following take-home message: either the number of observations $n$ is sufficiently small or the number of distinct values $k$ should be sufficiently large, and how large it must be depends on the choice of discount $\sigma$, namely $k\gtrsim n^{\frac{\sigma}{1+\sigma}}$.
@@
Optimal transport between random measures. The idea of endowing a space of measures with a distance and using it as a cost for an optimal transport distance on the laws of random measures is not new. This has been done between laws of random probability measures when the distance is an optimal transport distance, with recent applications in machine learning (Yurochkin et al., 2019, Dukler et al., 2019) and Bayesian statistics (Nguyen, 2016, Catalano and Lavenant, 2024). In a different direction, Huesmann (2016) extended this “OT over OT” problem to random measures that are equivariant, which implies that their law is invariant under translation. This corresponds to measures with infinite mass (because the space is unbounded), and the average transport cost is used as a measure of distance. Erbar et al. (2025) then studied the gradient flows in the geometry generated by this optimal transport problem. There have also been proposals of distances tailored for laws of point processes; see, e.g., Barbour and Brown (1992), Schuhmacher and Xia (2008), Decreusefond et al. (2016), and references therein. However, we focus on general random measures, not only point processes, which forces us to consider also the distribution of the jumps of the random measures, and not only the one of their atoms.
Optimal transport on CRMs. The first proposals to define a distance between CRMs with tractable upper bounds can be found in Catalano et al. (2020, 2021), where the authors propose using
@@
which is a distance between CRMs (Catalano et al., 2024) and leads to closed-form expressions whenever i) the CRMs do not have fixed atoms; ii) the CRMs are almost surely (a.s.) finite and homogeneous, i.e., $\nu^{i}=\rho^{i}\otimes P^{i}$ is a product measure; iii) the CRMs share the same base measure, i.e., $P^{1}=P^{2}$. These conditions are usually met in the setting of Catalano et al. (2024) but are very restrictive in our framework since i) and ii) typically hold a priori but not a posteriori, and iii) entails that the atoms of the CRM have the same distribution, which is likely to be the first feature to change when choosing a different prior. Therefore, we need to define the adapted extended Wasserstein discrepancy, a new OT distance on Lévy intensities with same mean measure that considers the joint distribution of the atoms and the jumps instead. Remarkably, it coincides with (3) whenever conditions i), ii), and iii) apply (see Remark 2 below).
@@
Merging of opinions. The notion of merging of opinions was introduced and thoroughly studied in the acclaimed work of Blackwell and Dubins (1962), which holds for a very general setup beyond exchangeability. In principle, their results cover model (1). However, in this setting, their assumptions can be considered restrictive for at least two reasons. First of all the merging is guaranteed for prior laws on $(X_{n})_{n\geq 1}$ that are absolutely continuous with one another. Whereas this holds in most parametric settings, it is not usually the case in nonparametric ones, where non-dominated models are the norm. As a classical example, two Dirichlet processes with different total base measures are mutually singular (Korwar and Hollander, 1973). It has also been observed by Diaconis and Freedman (1986), who avoid this condition by assessing the merging under a weaker topology. Moreover, both Blackwell and Dubins (1962) and Diaconis and Freedman (1986) assume that the data-generating mechanism coincides with one of the models under consideration, which may well not be the case in practice. Our aim is to overcome both restrictions and build a framework that does not require any assumption on the data. Indeed, we propose one of the few asymptotic analyses that does not put strong assumptions on the distribution of the data (e.g., to be a sampled from the model or i.i.d. from a true distribution): this is possible because we focus on the similarities between specific versions of the posteriors rather than on the adherence of the learning mechanism to the unknown truth.
@@
Frequentist consistency and contraction rates. In the same spirit, our framework goes beyond frequentist consistency properties in that it does not assume the data to be i.i.d. from a true distribution. However, if one were willing to make this assumption, consistency of the model with respect to each prior would entail merging of opinions but would not provide meaningful information about the convergence rate. Indeed the contraction rate of each prior to the truth could be (and usually is) considerably slower than the merging rate. For example, in this work we show that two Dirichlet processes with different parameters always merge at the rate 1/$n$, independently of the dimension of the space. This is much faster than the parametric rate of $1/\sqrt{n}$, which is the best possible one for the contraction rate, and moreover we expect the latter to deteriorate when the dimension increases. A merging rate that is faster than the contraction rate for a class of priors may suggest that it is the empirical part of the posterior - rather than the prior - that is slowing down convergence. Thus, not only our framework has the advantage of not putting assumptions on the data-generating process, it also provides new perspectives when one is willing to put them.
@@
Sensitivity analysis and robustness. To our knowledge, this work provides the first theoretical setup of global sensitivity analysis in Bayesian nonparametric non-dominated models, in the sense that it studies the performance of the model of interest with respect to any choice of prior. Previous studies on sensitivity for BNP models usually consist in varying the parameters of the priors or taking functional perturbations (Nieto-Barajas and Prünster, 2009) and providing empirical comparisons between the posteriors. In a dominated setting, more classical distances between (posterior) densities can be considered, as in Saha and Kurtek (2019). A recent work of Giordano et al. (2022) focuses on local sensitivity, which studies robustness with respect to small perturbation of the prior through differential approximations (Gustafson, 1996) and can be very useful when posterior distributions are not known in closed form. Since local sensitivity focuses on small perturbations, a good impact is always a small impact. On the contrary, our framework does not limit itself to small perturbations, and thus a big impact of the prior can also be regarded as a benign sign of flexibility of the model. Although our statistical analysis focuses on the latter aspect, our results may be read or adapted to treat robustness as well.
@@
Structure of the paper. In Section 2 we define the Wasserstein over BL distance between the laws of random measures. We characterize its topology and provide a general lower bound, specializing it to Poisson random measures. In Section 3 we focus on CRMs and provide a universal upper bound in terms of the adapted extended Wasserstein discrepancy between their Lévy intensities. The bound is specialized to Poisson random measures and extended to Cox CRMs with a random Lévy intensity. In Section 4 we establish the class of (Cox) CRMs that are identifiable for model (1) with respect to normalization, leading to the notion of scaled (Cox) CRMs. The adapted extended Wasserstein discrepancy is a distance on this class: in Section 5 we study its finite sample and asymptotic behavior to investigate the merging of opinions for BNP models. All proofs are deferred to the Supplementary Material (Catalano and Lavenant, 2025). Table 1 summarizes the distances that we use in this work.
Notations for random quantities. We denote deterministic scalar quantities with lowercase letters (e.g., $c,\alpha$) and real-valued random variables with capital letters (e.g., $C,U$). Random objects on spaces of measures are distinguished by a $\sim$, e.g., $\tilde{\nu}$ denotes a random measure while $\nu$ a deterministic measure.
@@
$\mathcal{W}_{d_{\mathbb{X}}}$ 11-Wasserstein distance. $\mathcal{P}(\mathbb{X})$ Eq. (4)
$\mathcal{W}_{d_{\mathbb{X},\text{t}}}$ 11-Wasserstein distance with metric $d_{\mathbb{X},\text{t}}$. $\mathcal{P}(\mathbb{X})$ Eqs. (4) and (6)
@@ 2.1 Distances between measures
$\mathcal{W}_{d_{\mathbb{Y}}}(P^{1},P^{2})=\min_{\pi\in\Pi(P^{1},P^{2})}\mathbb{E}_{(X,Y)\sim\pi}(d_{\mathbb{Y}}(X,Y))=\sup_{f}\left\{\int_{\mathbb{Y}}f\,\mathrm{d}P^{2}-\int_{\mathbb{Y}}f\,\mathrm{d}P^{1}\ :\ f\text{ is }1\text{-Lipschitz}\right\},$ (4)
where the minimum is taken over $\Pi(P^{1},P^{2})$ the set of joint probability distributions having marginals $P^{1},P^{2}$, and the equality follows by Kantorovich duality. Here $1$-Lipschitz means $|f(x)-f(y)|\leq d_{\mathbb{Y}}(x,y)$ for all $x,y\in\mathbb{Y}$. In the sequel we will use this definition for $(\mathbb{Y},d_{\mathbb{Y}})=(\mathbb{X},d_{\mathbb{X}})$, but also when $\mathbb{Y}=\mathcal{M}_{B}(\mathbb{X})$ or $\mathbb{Y}=\mathcal{M}((0,+\infty)\times\mathbb{X})$ is a space of measures.
Restricting to the case $(\mathbb{Y},d_{\mathbb{Y}})=(\mathbb{X},d_{\mathbb{X}})$, note that with $\mathcal{W}_{d_{\mathbb{X}}}$ we cannot compare measures of different mass, that is, $\mathcal{W}_{d_{\mathbb{X}}}(P^{1},P^{2})$ becomes infinite if $P^{1}(\mathbb{X})\neq P^{2}(\mathbb{X})$. This can be seen by taking constant functions $f(x)=c$ and letting $c\to\pm\infty$. A natural generalization is the Bounded Lipschitz distance $\mathrm{BL}$, which replaces the supremum over $1$-Lipschitz functions in (4) by a supremum over $1$-bounded and $1$-Lipschitz functions. Specifically, we define
$\mathrm{BL}(\mu^{1},\mu^{2})=\sup_{f}\left\{\int_{\mathbb{X}}f\,\mathrm{d}\mu^{2}-\int_{\mathbb{X}}f\,\mathrm{d}\mu^{1}\ :\ f\text{ is }1\text{-bounded and }1\text{-Lipschitz}\right\},$ (5)
where $1$-bounded means $|f(x)|\leq 1$ for all $x\in\mathbb{X}$. The $\mathrm{BL}$ distance metrizes the weak convergence on $\mathcal{M}_{B}(\mathbb{X})$ (see, e.g., Theorem 8.3.2 in Bogachev (2007)). We note that some works call $\mathrm{BL}$ the Kantorovich-Rubinstein or flat metric and rather define the function class of the Bounded Lipschitz distance to be $\{f:\|f\|_{L}+\|f\|_{\infty}\leq 1\}$, where $\|\cdot\|_{L}$ is the Lipschitz norm. The induced distance, first defined in Fortet and Mourier (1953) to study the convergence of the empirical measure, is easily shown to be strongly equivalent to our definition (see, e.g., Section 8.3 of Bogachev (2007)).
Regarding the comparison between $\mathcal{W}_{d_{\mathbb{X}}}$ and $\mathrm{BL}$ on $\mathcal{P}(\mathbb{X})$, note that we always have $\mathrm{BL}\leq\mathcal{W}_{d_{\mathbb{X}}}$. However, the discrepancy can be large: $\mathrm{BL}$ is always bounded by $2$, while $\mathcal{W}_{d_{\mathbb{X}}}$ can be arbitrary very large if the first moments of the measures are large. To perform a finer comparison, we introduce the “truncated” distance
@@
Thus, $\mathcal{W}_{d_{\mathbb{X},\text{t}}}$ and $\mathrm{BL}$ are strongly equivalent metrics when restricted to $\mathcal{P}(\mathbb{X})$, but $\mathcal{W}_{d_{\mathbb{X},\text{t}}}$ easier to interpret as it corresponds to a classical Wasserstein distance, though with respect to a truncated distance.